1. The reaction A → B + C is zero order with respect to A. When [A]0 = 0.544 M, the reaction is 28 % complete at 91.5 min. Calculate the half-life for this reaction.

2. Consider the following reaction. Calculate the value of the rate constant..

2 A + B →→ 2 C

A B C Rate

Trial 1 0.64 0.56 0.012 7.463

Trial 2 1.28 0.56 0.012 14.926   

Trial 3 0.64 1.12 0.012 14.926

Trial 4 0.1 0.4    0.012 ?
Answer #1

we have:

[A]o = 0.544 M

28 % is completed. So, remaining is 72 %

[A] = 72 % of 0.544 M

= 72*0.544/100

= 0.392 M

t = 91.5 min

use integrated rate law for 0 order reaction

[A] = [A]o - k*t

0.392 = 0.544 - k*91.5

k = 1.661*10^-3 min M/min

Given:

k = 1.661*10^-3 M/min

use relation between rate constant and half life of 0 order reaction

t1/2 = [A]o/(2*k)

= 0.544/(2*1.661*10^-3)

= 1.637*10^2 min

Answer: 164 min

Only 1 question at a time please